# MMEvol: Empowering Multimodal Large Language Models with Evol-Instruct

### v2

@@ Abstract
The development of Multimodal Large Language Models (MLLMs) has seen significant advancements with increasing demands in various fields (e.g., multimodal agents, embodied intelligence). While model-driven approaches attempt to enhance MLLMs capabilities through diverse architectures, the gains have become increasingly marginal. Conversely, data-driven methods, which scale up image-text instruction data, are more effective but face limited data diversity and complexity challenges. The absence of high-quality data constitutes a significant development barrier for MLLMs. To address the data quality bottleneck, we propose MMEvol, a novel multimodal instruction data evolution framework. This framework iteratively improve data quality through a refined combination of fine-grained perception, cognitive reasoning, and interaction evolution, generating a more complex and diverse image-text instruction dataset that empowers MLLMs with enhanced capabilities. Beginning with an initial set of instructions, SEED-163K, we utilize MMEvol to systematically broaden the diversity of instruction types, extend visual reasoning steps to improve cognitive reasoning abilities, and thoroughly explore fine-grained information within images to enhance visual understanding and robustness. To comprehensively evaluate the effectiveness of our approach, we conduct extensive qualitative analysis and quantitative experiments across 13 vision-language tasks. Compared to baseline models trained with the initial seed data, the results demonstrate that our method achieves an average accuracy improvement of 3.1 percentage points. Furthermore, our approach reaches state-of-the-art (SOTA) performance in nine tasks using significantly less data compared to state-of-the-art models. The project page is available at https://rainbowluocs.github.io/MMEvolWeb.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) (Liu et al., 2024b; a; Li et al., 2023b; Dong et al., 2023; Sun et al., 2023b; Dai et al., 2024; Luo et al., 2024; Qi et al., 2024) have seen rapid development over the past two years and have become the preferred approach for various vision-language tasks (Kembhavi et al., 2016; Fu et al., 2024; Zhang et al., 2024a; Qian et al., 2024). By aligning visual encoders (Radford et al., 2021; Zhai et al., 2023; Sun et al., 2023a) with LLMs (Touvron et al., 2023; Bai et al., 2023; Lu et al., 2024; Young et al., 2024; Tao et al., 2024), and employing large-scale coarse-grained image-text pre-training (Zhu et al., 2024; Schuhmann et al., 2022; 2021) followed by small-scale instruction-tuning (Chen et al., 2024a; Liu et al., 2024b), MLLMs have demonstrated impressive capabilities across numerous vision-language tasks and are widely applied in many domains (e.g., multimodal agents, embodied intelligence). Model-driven approaches (Luo et al., 2024; Liu et al., 2024a; Tong et al., 2024; Zhang et al., 2024b) aim to integrate knowledge from images and text more efficiently by designing different network architectures to improve MLLMs performance. However, their effectiveness is diminishing due to the lack of high-quality data, and redundant model designs struggle to realize their potential and push the boundaries of model intelligence. In contrast, data-driven methods (Liu et al., 2024b; Chen et al., 2024a; Yu et al., 2023; Liu et al., 2024c; Fang et al., 2024; Chen et al., 2023), while more effective, often produce data with limited diversity and complexity. The lack of high-quality data remains a significant obstacle in the development of MLLMs. Therefore, the need of developing automated methods that can generate more challenging and diverse instructional data at a relatively low cost is urgent for empowering MLLMs with enhanced capabilities.
@@
Figure 3: Prefix-Prompt of MMEvol. The top block showcases the contexts such as caption and visual object locations, and the middle block demonstrates vision/la nguage-centered atomic propositions and evolution objective (described later). Additionally, we endow vision capabilities with pseudo-function calls to enhance visual reasoning during evolutionary processes. Finally, the bottom block further elucidates the organized seed sample, which is subsequently sent to the MLLM for rewriting.
@@ 2 Method
In this section, we first introduce the curation of seed instruction data and then elaborate on the methodological details of MMEvol. Due to the space limitation, we simplify the seed data curation process and prompt templates. More details can be found in the Appendix D.
@@ 2.2 Methodological Details
Figure 4: Fine-grained perceptual evolution prompt and data example. Fine-grained perceptual evolution can generate samples with more detailed visual information, enhancing data diversity, which are marked with different colors for better visualization.
@@ 3.1 Benchmarks
Skills Sources Skills Sources
VQA VQAv2 (Goyal et al., 2017) General Knowledge MMEC (Fu et al., 2023)
Knowledge Leakage MMStar (Chen et al., 2024b) MMMU (Yue et al., 2024)
Math Reasoning MathVistaM (Lu et al., 2023) Hallucination POPE (Li et al., 2023c)
OCR Related AI2D (Kembhavi et al., 2016) HallBench  (Guan et al., 2023)
@@ 3.2 Implementation Details
Figure 8: MMEvol continuously enhances instruction data complexity and diversity over evol-instruct. The sample is from SEED-163K. We mark fine-grained visual information in red, new instructions form in green, and longer reasoning steps in blue.
@@ 3.3 Qualitative Analysis
Additionally, we randomly sample 30K data points from the seed data and conduct qualitative analysis on the instruction data before and after evolution. As shown in Fig. 10, the evolved data is notably more complex. Specifically, each evolved instruction involves 0.68 more atomic abilities in Fig. 10(a) and has an average visual operation chain reasoning length of 0.86 longer compared with pre-evolution in Fig. 10(b). As we can see from Fig. 10(c), the average difficulty score of each evolution round increases progressively, demonstrating the effectiveness of cognitive reasoning evolution in increasing instruction data complexity.
@@ 3.4 Ablation Study
We conduct ablation studies on seven vision-language benchmarks to explore the effects of instruction evolution and elimination. As shown in Table 2, different evolution process can be orthogonally superimposed on each other to continuously enhances data diversity and complexity. leading to an average performance gain of 3.8 points across multiple vision-language benchmarks. However, the absence of instruction elimination introduces harmful data from failed evolutions, which inevitably reduces the model’s resistance to hallucinations by 1.2 points on POPE  (Li et al., 2023c). When both instruction evolution and instruction elimination are employed, instruction elimination filters out harmful data from failed evolutions, further improving the quality and density of evolved data and enhancing the model’s performance by 0.9 points on average, particularly improving resistance to hallucinations by 1.7 points, which aligns with our qualitative analysis results in Section 3.3.
Table 2: Ablation study on instruction evolution and instruction elimination. The application of instruction evolution alone enhances the complexity and diversity of the data, whereas the integration of instruction elimination further refines data quality, markedly reducing the occurrence of visual hallucinations.
@@ 3.5 Benchmark Comparison
In comparison to the open-source SOTA model MiniCPM-v2.5 (Yao et al., 2024), despite a considerable difference in training data volume, our model still delivers better results, particularly showing improvements in instruction following, visual hallucinations, and visual reasoning with gains of $\uparrow$3.1 points on HallBench, $\uparrow$2.5 points on MIA, and $\uparrow$13.6 points on MMSInst respectively. This demonstrates that our method enhances the model’s visual reasoning and instruction following, reduces visual hallucinations, and improves other general capabilities, consistent with our findings from ablation studies and qualitative analyses. By using our data and the leading large language model Qwen2, we can train a superior MLLM from scratch in only one day using 4×8 A100 GPUs, further validating that high-quality instruction data is more important than large-scale low-quality data.
Table 3: Comparison with state-of-the-art methods on 13 visual-language benchmarks. Our models consistently improve LLaVA-NeXT under a head-to-head comparison, using the same prompts and the same base LLM, showing the effectiveness of enhanced pretraining data quality. We mark the best performance bold and the second-best underlined. “PT” denotes pre-training data scale, and “IT” denotes instruction tuning data scale.
@@
Seed-8B 0.6M 1.1M 82.5 64.8 41.3 47.4 60.8 47.7 38.0 72.1 85.3 69.4 44.2 59.9 26.2 56.9
Evol-8B 0.6M 1.6M 83.4 65.0 47.8 50.1 62.3 50.0 40.8 73.9 86.8 78.8 46.4 62.6 32.3 60.0
Evol-Qwen2-7B 0.6M 1.6M 83.1 65.5 55.8 51.6 64.1 52.4 45.1 74.7 87.8 77.6 47.7 63.9 41.8 62.4
@@ 5 Conclusions
Limitations. Due to resource constraints, we only performed evolution on 163K samples and conducted experiments with an 8B scale model. Expanding the dataset and using larger-scale models could yield even better results. We plan to explore these avenues in future work.
@@
Table 4: The mixture of training recipe datasets with corresponding categories and sources.
@@
General Knowledge OKVQA (Marino et al., 2019), A-OKVQA (Schwenk et al., 2022) 243K 14.9%
@@
Visual Reasoning GQA (Hudson & Manning, 2019) 72K 4.4%
Visual Grounding RefCOCO (Kazemzadeh et al., 2014) 48K 2.9%
OCR Related OCR-VQA (Mishra et al., 2019), TextVQA (Singh et al., 2019) 270K 16.5%
@@
General Knowledge Design2Code (Si et al., 2024a), MathVision (Wang et al., 2024a) 85K 5.2%
@@
Appendix A Curation Details of Seed Data
LLaVA-Instruct (Liu et al., 2024b) is a dataset of image-text instructions based on the COCO (Chen et al., 2015) data source and generated using the OpenAI ChatGPT API. The image-text instruction format in this dataset primarily includes three types: dialogue-based question-answering, global descriptions, and complex reasoning. ShareGPT4V (Chen et al., 2023), on the other hand, is a dataset constructed or rewritten using the OpenAI GPT-4V API, based on image-text pairs from SAM (Kirillov et al., 2023), COCO, and other sources to introduce richer details into captions. Both LLaVA-Instruct and ShareGPT4V significantly advance the development of MLLMs (Hu et al., 2022; Li et al., 2023d; Si et al., 2024b) and are widely used. We integrate samples from these two datasets containing the same image by concatenating the corresponding instruction data lists. For samples with global descriptions but no instruction data, we use the GPT-4o-mini API to supplement the missing instruction data, similar to LLaVA-Instruct, resulting in a combined dataset of 133K samples. To ensure the diversity of the seed data, we also include additional scientific chart data. Specifically, we sample 30K entries from Cambrain-1 (Tong et al., 2024), covering various types of image-text instructions such as code generation, chart interpretation, scientific question-answering, document understanding, and mathematical reasoning, ultimately forming a seed dataset of 163K image-text instructions.
@@ Appendix C Additional Visualization Results
We also present additional visualization results to demonstrate the capabilities of our model. As shown in Fig. 12, our model trained on this data exhibits strong visual reasoning, instruction following, and fine-grained perception capabilities. Additionally, it identifies nuances in meme content, validating the effectiveness and efficiency of MMEvol.
Appendix D Complete Evolution Prompt Template

### v4

@@ Abstract
The development of Multimodal Large Language Models (MLLMs) has seen significant advancements with increasing demands in various fields (e.g., multimodal agents, embodied intelligence). While model-driven approaches attempt to enhance MLLMs capabilities through diverse architectures, the gains have become increasingly marginal. Conversely, data-driven methods, which scale up image-text instruction data, are more effective but face limited data diversity and complexity challenges. The absence of high-quality data constitutes a significant development barrier for MLLMs. To address the data quality bottleneck, we propose MMEvol, a novel multimodal instruction data evolution framework. This framework iteratively improve data quality through a refined combination of fine-grained perception, cognitive reasoning, and interaction evolution, generating a more complex and diverse image-text instruction dataset that empowers MLLMs with enhanced capabilities. Beginning with an initial set of instructions, SEED-163K, we utilize MMEvol to systematically broaden the diversity of instruction types, extend visual reasoning steps to improve cognitive reasoning abilities, and thoroughly explore fine-grained information within images to enhance visual understanding and robustness. To comprehensively evaluate the effectiveness of our approach, we conduct extensive qualitative analysis and quantitative experiments across 13 vision-language tasks. Compared to baseline models trained with the initial seed data, the results demonstrate that our method achieves an average accuracy improvement of 3.1 percentage points. Furthermore, our approach reaches state-of-the-art (SOTA) performance in nine tasks using significantly less data compared to state-of-the-art models. The project page is available at https://mmevol.github.io/ Our code, weights, and data are all publicly available at .
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) (Liu et al., 2024b; a; Li et al., 2023c; Dong et al., 2023; Sun et al., 2023b; Dai et al., 2024; Luo et al., 2024; Qi et al., 2024) have seen rapid development over the past two years and have become the preferred approach for various vision-language tasks (Kembhavi et al., 2016; Fu et al., 2024; Zhang et al., 2024a; Qian et al., 2024). By aligning visual encoders (Radford et al., 2021; Zhai et al., 2023; Sun et al., 2023a) with LLMs (Touvron et al., 2023; Bai et al., 2023; Lu et al., 2024; Young et al., 2024; Tao et al., 2024), and employing large-scale coarse-grained image-text pre-training (Zhu et al., 2024; Schuhmann et al., 2022; 2021) followed by small-scale instruction-tuning (Chen et al., 2024a; Liu et al., 2024b), MLLMs have demonstrated impressive capabilities across numerous vision-language tasks and are widely applied in many domains (e.g., multimodal agents, embodied intelligence). Model-driven approaches (Luo et al., 2024; Liu et al., 2024a; Tong et al., 2024; Zhang et al., 2024b) aim to integrate knowledge from images and text more efficiently by designing different network architectures to improve MLLMs performance. However, their effectiveness is diminishing due to the lack of high-quality data, and redundant model designs struggle to realize their potential and push the boundaries of model intelligence. In contrast, data-driven methods (Liu et al., 2024b; Chen et al., 2024a; Yu et al., 2023; Liu et al., 2024c; Fang et al., 2024; Chen et al., 2023), while more effective, often produce data with limited diversity and complexity. The lack of high-quality data remains a significant obstacle in the development of MLLMs. Therefore, the need of developing automated methods that can generate more challenging and diverse instructional data at a relatively low cost is urgent for empowering MLLMs with enhanced capabilities.
@@
Figure 3: Prompt Head of MMEvol. The top block showcases the contexts such as caption and visual object locations, and the middle block demonstrates vision/la nguage-centered atomic propositions and evolution objective (described later). Additionally, we endow vision capabilities with pseudo-function calls to enhance visual reasoning during evolutionary processes. Finally, the bottom block further elucidates the organized seed sample, which is subsequently sent to the MLLM for rewriting.
@@ 2 Method
In this section, we first introduce the curation of seed instruction data and then elaborate on the methodological details of MMEvol. Due to the space limitation, we simplify the seed data curation process and prompt templates. More details can be found in the Appendix E.
@@ 2.2 Methodological Details
Figure 4: Fine-grained perceptual evolution prompt and example. Fine-grained perceptual evolution can generate samples with more detailed visual information, enhancing data diversity, which are marked with different colors for better visualization.
@@ 3.3 Qualitative Analysis
We randomly sample 30K data points from the seed data and conduct qualitative analysis on the instruction data before and after evolution. As shown in Fig. 9, the evolved data is notably more complex. Specifically, each evolved instruction involves 0.68 more atomic abilities in Fig. 9(a) and has an average visual operation chain reasoning length of 0.86 longer compared with pre-evolution in Fig. 9(b). As we can see from Fig. 9(c), the average difficulty score of each evolution round increases progressively, demonstrating the effectiveness of cognitive reasoning evolution in increasing instruction data complexity.
@@ 3.4 Ablation Study
We conduct ablation studies on seven vision-language benchmarks to explore the effects of instruction evolution and elimination. As shown in Table 1, different evolution process can be orthogonally superimposed on each other to continuously enhances data diversity and complexity. leading to an average performance gain of 3.8 points across multiple vision-language benchmarks. However, the absence of instruction elimination introduces harmful data from failed evolutions, which inevitably reduces the model’s resistance to hallucinations by 1.2 points on POPE  (Li et al., 2023d). When both instruction evolution and instruction elimination are employed, instruction elimination filters out harmful data from failed evolutions, further improving the quality and density of evolved data and enhancing the model’s performance by 0.9 points on average, particularly improving resistance to hallucinations by 1.7 points, which aligns with our qualitative analysis results in Section 3.3.
In addition, we conduct an ablation study starting from 1k seed data, focusing on the ratios of three evolutionary directions. The results, shown in Table 2, indicate that equal probabilities for the three directions yield the best average performance, confirming their importance in enhancing the complexity and diversity of augmented data. To demonstrate the effectiveness of our prompt design with explicit complexity and diversity metrics, we replace the evolved prompts with their simplest basic versions and conduct experiments on the 1k data. As shown in Table 3, measurable complexity and diversity metrics significantly improve the quality of data evolution. Please refer to Appendix D for more ablation study.
Table 1: Ablation study on instruction evolution and instruction elimination (6K). The application of instruction evolution alone enhances the complexity and diversity of the data, whereas the integration of instruction elimination further refines data quality, markedly reducing the occurrence of visual hallucinations.
@@
Table 2: Ablation study on instruction evolution ratio. The highest average performance can be achieved when the ratios of the three evolutionary directions are equal, thereby demonstrating that all three directions are equally important for the diversity and complexity of the evolutionary instruction data.
FP-Evol I-Evol CR-Evol I-Elim MMStar MathVistaM MMEC AI2D HallBench MMMUV RWQA AVG.
2/3 1/6 1/6 36.9 26.3 31.0 54.0 44.8 34.4 51.4 39.8
1/6 2/3 1/6 34.3 25.4 29.2 53.2 43.5 35.8 52.6 39.2
1/6 1/6 2/3 36.3 26.7 32.5 54.3 44.0 35.2 51.1 40.0
1/3 1/3 1/3 37.9 26.1 31.3 55.1 43.8 35.8 53.2 40.5
Table 3: Ablation study on prompt version. Utilizing our meticulously designed prompts significantly enhances the diversity and complexity of the data, thereby making the evolutionary process more efficient. The symbol ✔indicates that during the evolutionary process, the prompt has been replaced with the base version.
FP-Evol I-Evol CR-Evol I-Elim MMStar MathVistaM MMEC AI2D HallBench MMMUV RWQA AVG.
34.7 25.7 29.9 54.1 42.1 35.5 49.8 38.8
35.7 25.9 30.3 54.8 42.9 35.2 51.2 39.4
36.5 25.4 30.8 55.0 43.6 35.4 52.4 39.9
37.9 26.1 31.3 55.1 43.8 35.8 53.2 40.5
@@ 3.5 Benchmark Comparison
In comparison to the open-source SOTA model MiniCPM-v2.5 (Yao et al., 2024), despite a considerable difference in training data volume, MMEvol-8 still delivers better results, particularly showing improvements in instruction following, visual hallucinations, and visual reasoning with gains of $\uparrow$3.1 points on HallBench, $\uparrow$2.5 points on MIA, and $\uparrow$13.6 points on MMSInst respectively. This demonstrates that our method enhances the model’s visual reasoning and instruction following, reduces visual hallucinations, and improves other general capabilities, consistent with our findings from ablation studies and qualitative analyses. By using our data and the leading large language model Qwen2, we can train a superior MLLM from scratch in only one day using 4×8 A100 GPUs, further validating that high-quality instruction data is more important than large-scale low-quality data.
Table 4: Comparison with state-of-the-art methods on 13 visual-language benchmarks. Our models consistently improve LLaVA-NeXT under a head-to-head comparison, using the same prompts and the same base LLM, showing the effectiveness of enhanced pretraining data quality. “PT” denotes pre-training data scale, “IT” denotes instruction tuning data scale and “*” denotes the baseline model trained on the seed dataset. We mark the best performance bold and the second-best underlined.
@@
LLaVA-NeXT*-8B 0.6M 1.1M 82.5 64.8 41.3 47.4 60.8 47.7 38.0 72.1 85.3 69.4 44.2 59.9 26.2 56.9
LLaVA-NeXT*-Qwen2-7B 0.6M 1.1M 82.5 64.9 44.6 48.9 61.7 49.3 41.7 73.3 86.4 70.2 44.7 61.0 30.1 58.4
MMEvol-8B 0.6M 1.6M 83.4 65.0 47.8 50.1 62.3 50.0 40.8 73.9 86.8 78.8 46.4 62.6 32.3 60.0
MMEvol-Qwen2-7B 0.6M 1.6M 83.1 65.5 55.8 51.6 64.1 52.4 45.1 74.7 87.8 77.6 47.7 63.9 41.8 62.4
@@ 5 Conclusions
6 Limitation
Due to resource limitation, we only performed evolution on 163K samples (approximately 12% in original data recipes) and conducted experiments with an 8B scale model. Expanding the dataset and using larger-scale models could yield even better results. We plan to explore these avenues in future work and replace the OpenAI GPT4o-mini API with open-sourced model like QWen2VL.
@@ 6 Limitation
Appendix A Curation Details of Seed Data
LLaVA-Instruct (Liu et al., 2024b) is a dataset of image-text instructions based on the COCO (Chen et al., 2015) data source and generated using the OpenAI ChatGPT API. The image-text instruction format in this dataset primarily includes three types: dialogue-based question-answering, global descriptions, and complex reasoning. ShareGPT4V (Chen et al., 2023), on the other hand, is a dataset constructed or rewritten using the OpenAI GPT-4V API, based on image-text pairs from SAM (Kirillov et al., 2023), COCO, and other sources to introduce richer details into captions. Both LLaVA-Instruct and ShareGPT4V significantly advance the development of MLLMs (Hu et al., 2022; Li et al., 2023e; Si et al., 2024b) and are widely used. We integrate samples from these two datasets containing the same image by concatenating the corresponding instruction data lists. For samples with global descriptions but no instruction data, we use the GPT-4o-mini API to supplement the missing instruction data, similar to LLaVA-Instruct, resulting in a combined dataset of 133K samples. To ensure the diversity of the seed data, we also include additional scientific chart data. Specifically, we sample 30K entries from Cambrain-1 (Tong et al., 2024), covering various types of image-text instructions such as code generation, chart interpretation, scientific question-answering, document understanding, and mathematical reasoning, ultimately forming a seed dataset of 163K image-text instructions.
Table 5: The mixture of training recipe datasets with corresponding categories and sources. We collect these public dataset form internet.
@@ Appendix A Curation Details of Seed Data
Knowledge OKVQA (Marino et al., 2019), A-OKVQA (Schwenk et al., 2022) 243K 14.9%
@@
Reasoning GQA (Hudson & Manning, 2019) 72K 4.4%
Grounding RefCOCO (Kazemzadeh et al., 2014) 48K 2.9%
OCR OCR-VQA (Mishra et al., 2019), TextVQA (Singh et al., 2019) 270K 16.5%
@@
Science/Code Design2Code (Si et al., 2024a), MathVision (Wang et al., 2024a) 85K 5.2%
@@ Appendix C Additional Visualization Results
Table 7: Human Evaluation of Instruction Evolution. The table shows the accuracy of manual evaluations by five experts on MMEvol’s evolution and filtering in three directions.
Data ID Expert Image Categories FP-Evol (0-5) I-Evol (0-5) CR-Evol (0-5) I-Elim (0-15)
0,1,3,4,5,6 0 LandMark, OCR, Human & Clothes, Traffic, Living room, Sport 5,4,4,5,5,4 5,4,3,4,5,4 5,3,4,5,4,4 15,13,13,14,13,14
7,8,9,10,11,12 1 Kitchen, Office supplies & Tools, Plants, Animal, Sport, LandMark 5,5,4,5,4,4 5,4,5,5,4,4 5,5,4,4,5,4 14,15,13,15,14,13
13,14,15,16,17,18 2 Foods, LandMark, OCR, Human & Clothes, Traffic, Sport 4,4,3,5,4,5 5,4,4,4,4,5 4,5,5,4,5,5 14,14,15,13,14,15
19,20,21,22,23,24 3 Foods, Sport, LandMark, Office supplies & Tools, Plants, Traffic 3,4,5,5,5,4 3,4,5,5,5,5 5,5,5,5,5,5 13,15,14,15,15,15
25,26,27,28,29,30 4 Animal, Sport, Traffic, Landmark, Sport, Office supplies & Tools 4,5,5,5,5,5 4,5,5,5,4,5 5,5,3,5,5,5 14,15,14,15,14,15
Average Scores 89.3% 88.7% 92% 94.5%
We plot the performance of the model at every 1k step across 9 evaluation datasets in Fig. 12 to observe the learning trends during training. We can observe that the model learns OCR-related capabilities and mathematical reasoning abilities relatively smoothly, while general perception and cognitive skills exhibit more challenges. This may stem from conflicts arising from multi-source training tasks. A phased learning approach based on the difficulty of different tasks could be adopted to achieve better performance. We also present additional visualization results to demonstrate the capabilities of our model. As shown in Fig. 13, our model trained on this data exhibits strong visual reasoning, instruction following, and fine-grained perception capabilities. Additionally, it identifies nuances in meme content, validating the effectiveness and efficiency of MMEvol.
To investigate the reliability of the rewrites produced by GPT-4-o-mini, we conducted a manual evaluation of the data before and after the evolution process. Specifically, we first extracted 30 images of various types from the seed data to ensure diversity, keeping 5 relevant question-answer pairs for each image. Subsequently, we carried out the corresponding evolution in three different directions, ultimately obtaining 450 evolved question-answer pairs, which were then subject to scoring and filtering. The results were distributed among five experts for manual evaluation of the accuracy of the model evolution and the scoring filter. The data is summarized in the Table 7. From the table, it is evident that the average success rate of evolution using MLLM can reach 90%, while the accuracy of the scoring filter can achieve 94%, indicating the reliability of MMEovel. Additionally, we provide detailed scoring cases in Fig. 15, highlighted in red.
Table 8: Ablation study on VLM for instruction evolution. If we utilize a more advanced open-source model such as Qwen2VL with 72 billion parameters, MMEvol could achieve further improvements.
VLM MMStar MathVistaM MMEC AI2D HallBench MMMUV RWQA AVG.
GPT4o-mini (3K) 37.9 26.1 31.3 55.1 43.8 35.8 53.2 40.5
Qwen2VL-72B (3K) 39.1 27.9 33.1 57.8 46.4 36.9 46.9 41.2
Appendix D Additional Ablation Study
As shown in the Table 8, we achieve better results by replacing the closed-source model with the more advanced open-source Qwen2-VL, demonstrating the scalability and promise of our method. Furthermore, to fully showcase our technological contribution, we use the state-of-the-art method MIMIC-IT Li et al. (2023b) to construct a dataset under the same seed data and API conditions, and conduct experiments. The results, as shown in the Table 9, indicate that our method significantly outperforms MIMIC-IT. Additionally, our method is not restricted by data format or limited task forms, making it more user-friendly. We also perform robustness testing on the initial data of our method by selecting samples with a complexity score below 5 to construct a low-quality 1k seed dataset and conduct three rounds of evolution. As shown in the Table 10, MMEvol is affected by low-quality initial seed samples, resulting in slightly lower outcomes compared to randomly sampled initial seed data. However, after multiple rounds of evolution, the gap narrows significantly, indicating MMEvol’s strong robustness and ability to quickly adapt and generate more complex and diverse high-quality samples.
Appendix E Complete Evolution Prompt Template